OpenStudy (anonymous):

If f'(x)=-f(x) and f(1)=1, then f(x)=

OpenStudy (anonymous):

Don't you know how to solve differential equations like this?

OpenStudy (anonymous):

No I do not. I am not good in Calculus.

OpenStudy (mathmate):

y'=-y with y(1)=1 Integrate both sides, and insert y(1)=1 to find the integration constant.

OpenStudy (mathmate):

Or it should be y'/y = -1 => dy/y = -dx, integrate both sides to get log(y) = -x+C => y=e^(-x)+C Now substitute y(1)=1 to find C and complete the solution.

OpenStudy (anonymous):

\[ y' = -y \\ \frac{dy}{dx} = -y \\ \frac{1}{y} \frac{dy}{dx} = -1 \\ \int \frac{1}{y} \frac{dy}{dx} dx = \int -1dx \\ \]Let \(u = y\) then \(du = \frac{dy}{dx} dx\) (from chain rule) \[ \int \frac{1}{u}du = -x +C_2 \\ \ln(u) + C_1 = -x + C_2 \\ \ln(y) = -x + C_3 \\ e^{\ln(y)} = e ^{-x + C_3} \\ y = C_4 e^{-x} \]Now we use \(y(1) = 1\) \[ 1 = C_4 e^{-1} \\ e = C_4 \]Plug back into original equation: \[ y = e e^{-x} \\ y = e^{1-x} \]

OpenStudy (mathmate):

Sorry, it should read: \( log(y)=−x+C=>y=Ce^{−x}\) Consequently substitute y(1)=1 to find C.